### v2

# Sums of four fourth power of primes

Abstract.
For any sufficiently large $\ell$, suppose that $\ell$ can be expressed as
$\ell=p_{1}^{4}+p_{2}^{4}+p_{3}^{4}+p_{4}^{4},$
where $p_{1},p_{2},p_{3},p_{4}$ are primes.For such $\ell$, in this paper we will use circle method and sieves to prove that the proportion of $\ell$ in positive integers is at least $\frac{1}{27241.64}$ .
Key words and phrases:
Waring–Goldbach problem, Circle method, Sieve methods
1. Introduction
For any positive integers $m$ and $k$, the Warning-Goldbach Problem is to discuss every positive number $\ell_{m,k}$ can be expressed as the sum of the mthm-th powers of $k$ primes $p_{1},p_{2},\cdots,p_{k}$, i.e.
(1.1) $\ell_{m,k}=p_{1}^{m}+p_{2}^{m}+\cdots p_{k-1}^{m}+p_{k}^{m}.$
When $m=1,k=2$, this is the famous Goldbach’s Conjecture, the difficult problem, which has not been successfully solved so far. But when $m=1$, $k=3$, the case of three prime numbers has been proven, which is the famous theorem of three prime numbers[9]. Furthermore, for higher-order terms like $m=3,k=4$, some scholars have already obtained some results. Let $\mathcal{L}_{3,4}$ be the set of integers $\ell_{m,k}$ that can be written as the formula (1.1) when $m=3,k=4$ . In 1949, Roth[5] showed that
$\sum_{\ell_{3,4}\leqslant N\atop\ell_{3,4}\in\mathcal{L}_{3,4}}1\gg\frac{N}{\log^{8}N}.$
Futhermore, from 2001 to 2003, Ren[2][3]improved Roth’s theorem to the extent that
$\sum\limits_{\begin{subarray}{c}\ell_{3,4}\leqslant N\\ \ell_{3,4}\in\mathcal{L}_{3,4}\end{subarray}}1\geqslant\beta_{3,4}N,$
where $\beta_{3,4}=1/320$. In addition, Liu[4] improved this result and obtained $\beta_{3,4}=1/173.12$ in2012.
The authors are interested in this problem especially for more cases $m$ and $k$. In this paper we consider the case $m=4,k=4$ and try to say something about it. Let $\ell$ be a positive integer, suppose $\ell$ can be expressed as the sum of the four powers of four prime numbers, i.e.
(1.2) $\ell=p_{1}^{4}+p_{2}^{4}+p_{3}^{4}+p_{4}^{4},$
define $\mathcal{L}$ as the set of all $\ell$, and assume that
(1.3) $\sum\limits_{\begin{subarray}{c}\ell_{\leqslant}N\\ \ell\in\mathcal{L}\end{subarray}}1\geqslant\beta N,$
Then we will consider the distribution of $\ell$. No one has yet conducted research on this problem, at least the authors have not see any relevant references. Therefore, we have the following theorem.
Theorem. $\beta=\frac{1}{27241.64}$ is acceptable in (1.3).
Note. Compared to the sums of powers of four cubes of primes, the density of sums of four fourth power of primes is much smaller, but if we expand the expression to the sums of more than four prime numbers, the density may become even higher, which is a problem worthy of further research.
2. Prerequisite Knowledge
Firstly, we introduce some definitions to prepare for the proof of the theorem.
Let NN be a large integer, $\delta$ be small enough and
$U=\left(\frac{N}{32(1+\delta)}\right)^{1/4},\quad V=U^{5/6}.$
Now, we define AA and $B$ as two sufficiently large integers, and $B$ is sufficiently large relative to AA. Define $\mathfrak{M}(q,a)$ as the interval $a/q-L^{B}/U^{4},a/q+L^{B}/U^{4}]$, then write $\mathfrak{M}$ for the union of all $\mathfrak{M}(q,a)$ with $1\leqslant a\leqslant q\leqslant L^{B}$ and $(a,q)=1.$ Obviously, $\mathfrak{M}(q,a)$ are disjoint. Define $\mathfrak{m}$ as the complement of $\mathfrak{M}$ in $L^{B}/U^{4},1+L^{B}/U^{4}]$.
For any real numbers $\lambda$ and $X$, define the following two integrals
$\Phi(\lambda,X)=\int\limits_{X}^{2X}e(u^{4}\lambda)\>du,\quad\Psi(\lambda,X)=\int\limits_{X}^{2X}\frac{e(u^{4}\lambda)}{\log u}\>du.$
Let
(2.1) $J(n)=\int_{-U^{4}L^{B}}^{U^{4}L^{B}}\Phi(\lambda,U)\overline{\Psi(\lambda,U)}\Big|\Psi(\lambda,U)\Big|^{2}\Big|\Psi(\lambda,V)\Big|^{4}e(-n\lambda)\>d\lambda.$
Define
(2.2) $\mathfrak{S}_{d}(n)=\sum_{q=1}^{\infty}T_{d}(n,q)\>,$
where
(2.3) $T_{d}(n,q)=\sum_{\overset{a=1}{(a,q)=1}}^{q}\frac{S(q,ad^{3})C^{3}(q,a)\overline{C^{4}(q,a)}}{q\varphi^{7}(q)}e\biggl(-\frac{an}{q}\biggr)$
and
(2.4) $S(q,a)=\sum_{m=1}^{q}e\biggl(\frac{am^{4}}{q}\biggr),\quad C(q,a)=\sum_{(m,q)=1}^{q}e\biggl(\frac{am^{4}}{q}\biggr).$
3. Some Lemmas
To prove the validity of the theorem, we also introduce the following lemmas:
Lemma 1. Let $0\leqslant|n|\leqslant N$, for each $m$ with $U<m\leqslant 2U$, denote by $R(m)$ the number of solutions of
$n=m^{4}+p_{2}^{4}+\cdots+p_{4}^{4}-p_{5}^{4}-\cdots-p_{8}^{4}$
with
$p_{2},p_{5},p_{6}\sim U,\quad p_{3},p_{4},p_{7},p_{8}\sim V.$
For $0<\xi<1/16$ and $D=N^{\xi}$, take $\mathfrak{S}_{d}(n)$ and $J(n)$ from the above (2.1), (2.2), define $E_{d}(n)$ as follows:
$\sum_{\begin{subarray}{c}m\sim U\\ m\equiv 0({\mathrm{mod}}~d)\end{subarray}}R(m)=\frac{\mathfrak{S}_{d}(n)}{d}J(n)+E_{d}(n).$
Then we have:
[…]
(ii) $J(n)$ is positive and satisfies
$J(n)\leqslant 11606V^{4}L^{-7}.$
(iii) For any complex numbers $\eta_{d}$ with $|\eta_{d}|\leqslant\tau(d)$, we hvae
$\sum_{d\leqslant D}\eta_{d}E_{d}(n)\ll V^{4}L^{-A}.$
Proof. Firstly, we prove (ii) of Lemma 1. From elementary estimation
$\int_{X^{4}}^{16X^{4}}e(\lambda u)\>du\leqslant\min(X^{4},|\lambda|^{-1})$
and integration by parts, we have
$\Phi(\lambda,X)=\frac{1}{4}\int_{X^{4}}^{16X^{4}}u^{-3/4}e(\lambda u)\,du\ll X^{-3}\min(X^{4},|\lambda|^{-1})$
and
$\Psi(\lambda,X)\ll X^{-3}\log^{-1}X\min(X^{4},|\lambda|^{-1}).$
We hope to calculate (2.1) by integrating the entire real axis, but this will result in the following error
$\ll\int_{U^{-4}L^{B}}^{\infty}|\Phi(\lambda,U)||\Psi(\lambda,U)|^{3}|\Psi(\lambda,V)|^{4}\>d\lambda$
$\ll U^{-12}V^{-12}L^{-7}\int_{U^{-4}L^{B}}^{\infty}\min(U^{4},|\lambda|^{-1})^{4}\min(V^{4},|\lambda|^{-1})^{4}\>d\lambda$
$\ll V^{4}L^{-3B}.$
By integral transformation, we have
$\int_{-\infty}^{\infty}\Phi(\lambda,U)\overline{\Psi(\lambda,U)}\big|\Psi(\lambda,U)\big|^{2}\big|\Psi(\lambda,V)\big|^{4}e(-n\lambda)\>d\lambda$
$=$ $\frac{1}{4}\int_{\mathcal{D}}\frac{d\nu_{1}\cdots d\nu_{4}du_{1}\cdots du_{3}}{\nu_{1}^{3/4}\cdots\nu_{4}^{3/4}u_{1}^{3/4}\cdots u_{4}^{3/4}\log\nu_{1}\cdots\log\nu_{4}\log u_{1}\cdots\log u_{3}},$
where
$\mathcal{D}=\left\{(v_{1},\ldots,v_{4},u_{1},\ldots,u_{3})\colon V^{4}\leqslant v_{1},\ldots,v_{4}\leqslant 16V^{4},\ U^{4}\leqslant u_{1},\ldots,u_{4}\leqslant 16U^{4}\right\}$
and $u_{4}=n+v_{1}+v_{2}-v_{3}-v_{4}+u_{1}+u_{2}-u_{3}.$
Then we get
(3.1) $\begin{split}J(n)=&\int_{-\infty}^{\infty}\Phi(\lambda,U)\overline{\Psi(\lambda,U)}|\Psi(\lambda,U)|^{2}|\Psi(\lambda,V)|^{4}e(-n\lambda)d\lambda\\ &+\>O(V^{4}L^{-3B})\\ =&\frac{1}{4}\int_{V^{4}}^{16V^{4}}\>\frac{dv_{1}}{v_{1}^{3/4}\log v_{1}}\cdots\int_{V^{4}}^{16V^{4}}\>\frac{dv_{4}}{v_{4}^{3/4}\log v_{4}}\int_{U^{4}}^{16U^{4}}\>\frac{du_{1}}{u_{1}^{3/4}\log u_{1}}\\ &\times\int_{U^{4}}^{16U^{4}}\frac{du_{2}}{u_{2}^{3/4}\log u_{2}}\int_{\max(U^{4},x-16U^{4})}^{\min(16U^{4},x-U^{4})}\frac{du_{3}}{u_{3}^{3/4}(x-u_{3})^{3/4}\log u_{3}}\\ &+\>O(V^{4}L^{-3B}).\end{split}$
Let the last integral in the equation be denoted as II, where $x=n+v_{1}+v_{2}-v_{3}-v_{4}+u_{1}+u_{2}$, now we calculate the upper bound of II. Firstly, regarding $x\leqslant 2U^{4}$ or $x\geqslant 32U^{4}$, in both cases, the integral region does not exist, so we have $2U^{4}<x<32U^{4}$. In this case, we record as $u_{3}=xu$, then there is
$\begin{split}I&\leqslant(1+\varepsilon)x^{-1/2}L^{-1}\int_{\max(U^{4}/x,1-16U^{4}/x)}^{\min(16U^{4}/x,1-U^{4}/x)}u^{-3/4}(1-u)^{-3/4}\,du\\ &\leqslant\frac{1+\varepsilon}{\sqrt{2}}U^{-2}L^{-1}\int_{1/17}^{16/17}u^{-3/4}(1-u)^{-3/4}\,du\\ &\leqslant\frac{1+\varepsilon}{\sqrt{2}}U^{-2}L^{-1}I^{*},\end{split}$
where we record as $I^{*}=\int_{1/17}^{16/17}u^{-3/4}(1-u)^{-3/4}\,du$, then substituting it into (3.1) yields the following
$J(n)$ $\leq\frac{1+\varepsilon}{4\sqrt{2}}I^{*}U^{-2}L^{-1}\Bigg(\>\int_{U^{4}}^{16U^{4}}\>\frac{du}{u^{3/4}\log u}\Bigg)^{2}\Bigg(\>\int_{V^{4}}^{16V^{4}}\>\frac{dv}{v^{3/4}\log v}\Bigg)^{4}+O(V^{4}L^{-3B})$
$<\left(\frac{24}{5}\right)^{4}\frac{4}{\sqrt{2}}(1+\varepsilon)I^{*}V^{4}L^{-7}+O(V^{4}L^{-3B})$
$<11606V^{4}L^{-7},$
where we have used the estimate $I^{*}<7.73$. This proves (ii). For the proofs of (i) and (iii), we can refer to the methods of Brüdern [1] and Ren[2]. This proves Lemma 1.
Lemma 2. For $(d,6)=1$, we get
$\begin{split}\mathfrak{S}_{d}(n)=&\{1+T_{1}(n,2)+T_{1}(n,2^{2})+T_{1}(n,2^{3})+T_{1}(n,2^{4})\}\\ &\times\prod_{\begin{array}[]{c}p\nmid d\\ p\neq 2\end{array}}\{1+T_{1}(n,p)\}\prod_{p\mid d}\{1+T_{p}(n,p)\}.\end{split}$
Proof. According to the definition in (2.3), it can be obtained through a simple proof $\mathfrak{S}_{d}(n)$ which is an integral function, then there is
$\begin{split}\mathfrak{S}_{d}(n)&=\prod_{p}\{1+T_{p}(n,p)+T_{p}(n,p^{2})+\cdots\}\\ &=\prod_{p\nmid d}\{1+T_{1}(n,p)+T_{1}(n,p^{2})+\cdots\}\\ &\times\prod_{p\mid d}\{1+T_{p}(n,p)+T_{p}(n,p^{2})+\cdots\}.\end{split}$
For $p\nmid d$, we have
$\begin{split}T_{d}(n,p)&=\sum_{\overset{a=1}{(a,p)=1}}^{p}\frac{S(p,ad^{3})C^{3}(p,a)\overline{C^{4}(p,a)}}{p\varphi^{7}(p)}e\left(-\frac{an}{p}\right)\\ &=\sum_{\overset{a=1}{(a,p)=1}}^{p}\frac{S(p,a)C^{3}(p,a)\overline{C^{4}(p,a)}}{p\varphi^{7}(p)}e\left(-\frac{an}{p}\right)=T_{1}(n,p).\end{split}$
According to Lemma 4 in Hua[6], $C(p^{t},a)=0$($p\neq 2$ and $t\geq 2$ or $p=2,t\geq 5$), substituting back to the original equation yields the Lemma 2.
Lemma 3.Define $K(n,p)$ as the number of solutions to the following equation
$y_{2}^{4}+\cdots+y_{4}^{4}-y_{5}^{4}-\cdots-y_{8}^{4}\equiv n(\text{mod}~p)$
where $1\leq y_{i}\leq p-1$ $(1\leq i\leq 8)$, then we get
$pK(n,p)=(p-1)^{7}+E,$
where
$E=\begin{cases}\>-(p-1),&p\mid n,\\ \>1,&p\nmid n,\end{cases}~~~~~~~~~~~~~~\textit{if}~p\equiv 3({\mathrm{mod}}~4);$
and
$|E|\leqslant(3\sqrt{p}+1)^{5}(p-1)(3p+1),~~~\textit{if}~p\equiv 1({\mathrm{mod}}~4).$
It follows that $K(n,p)>0$ for $p\geq 13$.
Proof. From the definition of $K(n,p)$, we have
$pK(n,p)=\sum_{a=1}^{p}C^{3}(p,a)\overline{C^{4}(p,a)}e\left(-\frac{an}{p}\right)=(p-1)^{7}+E,$
where
$E=\sum_{a=1}^{p-1}C^{3}(p,a)\overline{C^{4}(p,a)}e\biggl(-\frac{an}{p}\biggr).$
When $p\equiv 3(\text{mod }4)$ and $(p,a)=1$, by Lemma 4.3 in Vaughan[7], we get $S(p,a)=0$ thus C(p,a)=1C(p,a)=-1, therefore
$E=-\sum_{a=1}^{p-1}e\left(-\frac{an}{p}\right)=\begin{cases}-(p-1),&p\mid n,\\ 1,&p\nmid n.\end{cases}$
and when $p\equiv 1(\text{mod }4)$, by Lemma 4.3 of Vaughan[7] again, we have $|C(p,a)|\leqslant 3\sqrt{p}+1$, and for
$\sum_{a=1}^{p-1}|C(p,a)|^{2}=\sum_{a=1}^{p}|C(p,a)|^{2}-(p-1)^{2}.$
Obiviously, $\sum_{a=1}^{p}|C(p,a)|^{2}$ can be expressed as $p$ times the number of solutions to equation $x^{4}\equiv y^{4}({\mathrm{mod~}}p)$, $1\leqslant x$, $y\leqslant p-1$. For $p\equiv 1({\mathrm{mod~}}4)$,
$\sum_{a=1}^{p}|C(p,a)|^{2}=4p(p-1),$
then
$\sum_{a=1}^{p-1}|C(p,a)|^{2}=(3p+1)(p-1),$
thus
$\sum_{a=1}^{p-1}|C^{3}(p,a)\overline{C^{4}(p,a)}|\leqslant(3\sqrt{p}+1)^{5}(p-1)(3p+1).$
A simple calculation shows that $K(n,p)>0$ for $p\geq 7$, $p\equiv 1(\text{mod }4)$and for all $p\equiv 3(\text{mod }4)$, therefore we have chosen $p\geq 13$. This proves Lemma 3.
From the similar method of Lemma 3, we can get Lemma 4.
Lemma 4. For $i=1,2$, let $H(n,p^{i})$ denote the number of solution of
$x^{3}+y_{2}^{3}+\cdots+y_{4}^{3}-y_{5}^{3}-\cdots-y_{8}^{3}\equiv n\pmod{p^{i}}$
where $1\leq x\leq p^{i},1\leq y_{j}<p^{i}$ and $(y_{j},p)=1$. Thus
$pH(n,p)=p(p-1)^{7}+E^{*},$
where
$E^{*}=0,~~~~~\textit{ if}~p\equiv 3(\text{mod }4);$
and
$|E^{*}|\leq 3\sqrt{p}(3\sqrt{p}+1)^{5}(p-1)(3p+1),~~\textit{if}~p\equiv 1(\text{mod }4).$
It follows that $H(n,p)>0$ for all pp.
Lemma 5. Definne the functions
$\phi_{0}(u)=\frac{2e^{\gamma_{0}}}{u}\mathrm{log}(u-1)\quad\text{and}\quad\phi_{1}(u)=\frac{2e^{\gamma_{0}}}{u},$
where $2\leq u\leq 3$. Suppose $\omega(d)$ is a multiplicative function of $d$ satisfying the conditions
$0\leqslant\omega(p)<p\quad\text{and}\quad\omega(p^{l})=1+O(p^{-1}),$
for each prime number $p$ and natural number $l$. Let $X$ be a real number with $X>3$, for $r(x)$ be a non-negative arithmetical function, we define
$E_{d}=\sum_{\begin{smallmatrix}P<x<2P\\ x\equiv 0\pmod{d}\end{smallmatrix}}r(x)-\frac{\omega(d)}{d}X.$
Let $U,V$ and $z$ be positive real parameters satisfying the inequality
$2\leqslant\frac{\log(UV)}{\log z}\leqslant 3.$
For any sequences $\{a_{m}\}$ and $\{b_{k}\}$ with
$|a_{m}|\leqslant 1\quad and\quad|b_{k}|\leqslant 1,$
one has
$\sum_{1\leq m\leq U}a_{m}\sum_{1\leq k\leq V}b_{k}E_{mk}\ll X(\log X)^{-2}.$
Then, we write
$W(z)=\prod_{p<z}\>(1-\omega(p)/p),$
one has the lower bound
$\sum_{\begin{subarray}{c}P\leq x<2P\\ (x,\Pi(z))=1\end{subarray}}r(x)>XW(z)\left(\phi_{0}\left(\frac{\log(UV)}{\log z}\right)+O\left((\log\log X)^{-1/50}\right)\right),$
and also the upper bound
$\sum_{\begin{subarray}{c}P\leq x<2P\\ (x,\Pi(z))=1\end{subarray}}r(x)<XW(z)\left(\phi_{1}\left(\frac{\log(UV)}{\log z}\right)+O\left((\log\log X)^{-1/50}\right)\right).$
Proof. See [8], Lemma 9.1.
Lemma 6. (Mertens’ theorem)For prime number $p$, $x\geq e$, positive integers $k$, $l$ with $(k,l)=1$ and $k\leq\ln^{A}{x}$ for any positive integer A, then
$\prod_{p\leq x,p\equiv l({\mathrm{mod~}k})}\left(1-\frac{1}{p}\right)=\frac{e^{-\frac{1}{\varphi(k)}\left[\gamma+\ln\frac{\varphi(k)}{k}\right]}}{\ln^{\frac{1}{\varphi(k)}}x}\left\{1+O\left(e^{-c\ln^{\frac{3}{5}}x}\right)\right\}$
where $c$ is a positive absolute constant, $\gamma$ is the Euler’s constant, and the constant in $O$ is independent of $x$.
Lemma 7. For $N/9<\ell\leqslant N$, define $r(\ell)$ as the number of $\ell$ can be expressed in the form of (1.2) with
$p_{1},p_{2}\sim U,\qquad p_{3},p_{4}\sim V\,.$
Then we have
$\sum_{N/9<\ell\leqslant N}r^{2}(\ell)<bV^{4}L^{-8},$
where $b=3702014554$.
Proof. Denote $\mathcal{B}$ as the set of all prime numbers greater than 13, and define
$P(z)=\prod_{\begin{subarray}{c}p<z\\ p\in\mathcal{B}\end{subarray}}p.$
To prove Lemma 7, we hope to obtain an appropriate upper bound for the following equation
$\sum_{m\sim U\atop(m,P(z))=1}R(m).$
For this, we apply Lemma 5, there is no prime divisor beyond $\mathcal{B}$ for $d$, then we define
$\omega(d)=\mathfrak{S}_{d}(n)/\mathfrak{S}_{1}(n).$
Especially by Lemma 2, for $p\in\mathcal{B}$ we have
$\omega(p)=\frac{1+T_{p}(n,p)}{1+T_{1}(n,p)}.$
By (2.3), we get
$1+T_{p}(n,p)=\sum_{a=1}^{p}\frac{\overline{C(p,a)}|C(p,a)|^{6}}{(p-1)^{7}}e\Bigg(-\frac{an}{p}\Bigg)=p\frac{K(n,p)}{(p-1)^{7}}.$
hence
$\omega(p)=\frac{K(n,p)}{H(n,p)}.$
By Lemma 3 and Lemma 4, through simple calculations, for all $p\in\mathcal{B}$ and the positive integer $l$, we have
$0\leqslant\omega(p)<p,\quad\omega(p^{l})=1+O(p^{-1}).$
Now let $X=\mathfrak{S}_{1}(n)J(n)$, then by Lemma 1, we have
$\sum_{\begin{subarray}{c}m\sim U\\ m\equiv 0\mod d\end{subarray}}R(m)=\frac{\omega(d)}{d}X+E_{d}(n).$
Suppose $U_{0}\geqslant 1,\ V_{0}\geqslant 1,\ U_{0}V_{0}=D=N^{(1-\epsilon)/16}$ and $z=D^{1/2}$, then
$\frac{\log(U_{0}V_{0})}{\log z}=2.$
For any sequence of numbers $\{a_{m}\}$, $\{b_{k}\}$ with
$|a_{m}|\leq 1~~,~~|b_{k}|\leq 1,$
by Lemma 1, we have
$\sum_{1\leqslant m\leqslant U}a_{m}\sum_{1\leqslant k\leqslant V}b_{k}E_{mk}\ll\sum_{1\leqslant d\leqslant D}\tau(d)E_{d}\ll V^{4}L^{-A}.$
By Lemma 5, we get
$\sum\limits_{(m,P(z))=1}R(m)<e^{\gamma}(1+\varepsilon)J(n)\mathfrak{S}_{1}(n)W(z),$
where $\gamma$ denotes Euler’s constant, and
$W(z)=\prod_{\stackrel{{\scriptstyle p\in\mathcal{B}}}{{p<z}}}\biggl(1-\frac{\omega(p)}{p}\biggr).$
Actually, we have estimated the upper bound of $J(n)$ in Lemma 1. Next, we estimate the remaining part,
(3.2) $\begin{split}&\mathfrak{S}_{1}(n)W(z)\\ =&\{1+T_{1}(n,2)+T_{1}(n,2^{2})+T_{1}(n,2^{3})+T_{1}(n,2^{4})\}\\ &\times(1+T_{1}(n,3))(1+T_{1}(n,5))(1+T_{1}(n,7))(1+T_{1}(n,11))\\ &\times\prod_{13\leqslant p<N^{(1-\varepsilon)/32}}(1+T_{1}(n,p))\biggl(1-\frac{K(n,p)}{H(n,p)}\biggr)\\ &\times\prod_{p\geqslant N^{(1-\varepsilon)/32}}(1+T_{1}(n,p))\\ =&\{1+T_{1}(n,2)+T_{1}(n,2^{2})+T_{1}(n,2^{3})+T_{1}(n,2^{4})\}\\ &\times(1+T_{1}(n,3))(1+T_{1}(n,5))(1+T_{1}(n,7))(1+T_{1}(n,11))\\ &\times\prod_{13\leqslant p<N^{(1-\varepsilon)/32}}\left(1-\frac{1}{p}\right)\left(1+\frac{E^{*}-E}{\left(p-1\right)^{8}}\right)\\ &\times\prod_{p\geqslant N^{(1-\varepsilon)/32}}\left(1+\frac{E^{*}}{p(p-1)^{7}}\right).\end{split}$
For $n$ in the definition of $R(m)$, define $\rho(n)$ as the number of solutions to the following equation
$n=p_{1}^{4}+\cdots+p_{4}^{4}-p_{5}^{4}-\cdots-p_{8}^{4},$
where
$p_{1},p_{2},p_{5},p_{6}\sim U,\quad p_{3},p_{4},p_{7},p_{8}\sim V.$
We can naturally obtain
$\sum_{N/9<\ell\leqslant N}r^{2}(\ell)\leqslant\rho(0)$
and
$\rho(n)\leq\sum_{m\sim U\atop(m,P(z))=1}R(m).$
Now we are calculating the various parts in (3.2). Since we are actually just trying to obtain a suitable upper bound for $\rho(0)$, thus we assume $n=0$ in the following text.
According to the definition (2.3), (2.4), by simple calculation we have $S(2,1)=0$, then
$T_{1}(n,2)=\sum_{\overset{a=1}{(a,2)=1}}^{2}\frac{S(2,a)C^{3}(2,a)\overline{C^{4}(2,a)}}{2\varphi^{7}(2)}e\biggl(-\frac{an}{2}\biggr)=0.$
For $q=4$, we have
$T_{1}(n,4)=\sum_{\overset{a=1}{(a,4)=1}}^{4}\frac{S(4,a)C^{3}(4,a)\overline{C^{4}(4,a)}}{4\varphi^{7}(4)}e\biggl(-\frac{an}{4}\biggr),$
where
$S(4,a)=\sum_{m=1}^{4}e\biggl(\frac{am^{4}}{4}\biggr)=2+2e\biggl(\frac{a}{4}\biggl),C(4,a)=\sum_{(m,4)=1}^{4}e\biggl(\frac{am^{4}}{4}\biggr)=2e\biggl(\frac{a}{4}\biggl).$
Thus
$\begin{split}T_{1}(n,4)&=\sum_{\overset{a=1}{(a,4)=1}}^{4}\frac{(2+2e(\frac{a}{4}))(2e(\frac{a}{4}))^{3}\overline{2e(\frac{a}{4})}^{4}}{4\varphi^{7}(4)}e\biggl(-\frac{a\times 0}{4}\biggr)\\ &=\frac{1}{2}\sum_{\overset{a=1}{(a,4)=1}}^{4}(1+e(-\frac{a}{4}))=\frac{1}{2}(2+2\cos{\frac{\pi}{2}})=1.\end{split}$
Similarly, we have
$T_{1}(n,8)=\frac{1}{2}(4+2\cos{\frac{\pi}{4}}+2\cos{\frac{3\pi}{4}})=2$
and
$T_{1}(n,16)=\frac{1}{2}(8+2\cos{\frac{\pi}{8}}+2\cos{\frac{3\pi}{8}}++2\cos{\frac{5\pi}{8}}+2\cos{\frac{7\pi}{8}})=4.$
Thus
$1+T_{1}(n,2)+T_{1}(n,2^{2})+T_{1}(n,2^{3})+T_{1}(n,2^{4})=8.$
Also, for $q=3$, we have
$T_{1}(n,3)=\sum_{\overset{a=1}{(a,3)=1}}^{3}\frac{S(3,a)C^{3}(3,a)\overline{C^{4}(3,a)}}{3\varphi^{7}(3)}e\biggl(-\frac{an}{3}\biggr),$
where
$S(3,a)=\sum_{m=1}^{3}e\biggl(\frac{am^{4}}{3}\biggr)=1+2e(\frac{a}{3}),C(4,a)=\sum_{(m,4)=1}^{4}e\biggl(\frac{am^{4}}{4}\biggr)=2e(\frac{a}{3}).$
then
$\begin{split}T_{1}(n,3)&=\sum_{\overset{a=1}{(a,3)=1}}^{3}\frac{(1+2e(\frac{a}{3}))(2e(\frac{a}{3}))^{3}\overline{(2e(\frac{a}{3}))^{4}}}{3\varphi^{7}(3)}e\biggl(-\frac{a\times 0}{3}\biggr)\\ &=\frac{1}{3}(4+e(-\frac{1}{3})+e(-\frac{2}{3}))=\frac{1}{3}(4+2\cos{\frac{2\pi}{3}})=1.\end{split}$
Likewise, by a series of calculations we have $T_{1}(n,5)=\frac{1}{5}(16+2\cos{\frac{2\pi}{5}}+2\cos{\frac{4\pi}{5}})=\frac{1}{5}(16+\frac{\sqrt{5}-1}{2}-\frac{\sqrt{5}+1}{2})=3$, $T_{1}(n,7)=\frac{8}{729}$ and $T_{1}(n,11)=\frac{324}{171875}.$
Furthermore, for
$\prod_{13\leqslant p<N^{(1-\varepsilon)/32}}\left(1+\frac{E^{*}-E}{\left(p-1\right)^{8}}\right),$
by Lemma 3 and Lemma 4, for $p\equiv 3(\text{mod }4)$ we have
$|E-E^{*}|\leq p-1,$
then
$\begin{split}\prod_{\overset{13\leqslant p<N^{(1-\varepsilon)/32}}{p\equiv 3(\text{mod}4)}}\left(1+\frac{E^{*}-E}{\left(p-1\right)^{8}}\right)&\leq\prod_{\overset{p\geqslant 13}{p\equiv 3(\mathrm{mod~4})}}\left(1+\frac{1}{\left(p-1\right)^{7}}\right)\\ &<1\>+\>\sum_{n\geqslant 12}\>\frac{1}{n^{7}}<1\>+\>12^{-6},\end{split}$
and for $p\equiv 1(\text{mod }4)$, we have
$|E-E^{*}|\leq(3\sqrt{p}+1)^{6}(p-1)(3p+1).$
then for $p\geq 101$, by $1+x\leq e^{x}$, we have
$\prod_{\overset{p\geq 101}{p\equiv 1(\text{mod}4)}}\left(1+\frac{E^{*}-E}{\left(p-1\right)^{8}}\right)\leq\prod_{\overset{p\geq 101}{p\equiv 1(\text{mod}4)}}e^{\frac{E^{*}-E}{\left(p-1\right)^{8}}}=e^{\sum_{p\geq 101}{\frac{E^{*}-E}{\left(p-1\right)^{8}}}},$
where
$\sum_{p\geq 101}{\frac{|E^{*}-E|}{\left(p-1\right)^{8}}}\leq 0.305,$
thus
$\prod_{\overset{p\geq 101}{p\equiv 1(\text{mod}4)}}\left(1+\frac{E^{*}-E}{\left(p-1\right)^{8}}\right)\leq 1.357.$
In addition,
$\prod_{\overset{13\leq p\leq 101}{p\equiv 1(\text{mod}4)}}\left(1+\frac{E^{*}-E}{\left(p-1\right)^{8}}\right)\leq 11.773.$
Therefore we have
$\prod_{13\leqslant p<N^{(1-\varepsilon)/32}}\left(1+\frac{E^{*}-E}{\left(p-1\right)^{8}}\right)<15.976$
Furthermore, by Lemma 6, we have the estimate
$\prod_{13\leqslant p<N^{(1-\varepsilon)/32}}\left(1-\frac{1}{p}\right)<308e^{-\gamma}(1+\varepsilon)L^{-1}.$
It is not difficult to find that
$\frac{|E^{*}|}{p(p-1)^{7}}<\frac{1}{p^{2}}$
for $p$ large enough. Hence
$\prod_{p\geqslant N^{(1-\varepsilon)/32}}\left(1+\frac{E^{*}}{p(p-1)^{7}}\right)\leqslant\prod_{p\geqslant N^{(1-\varepsilon)/32}}\left(1+\frac{1}{p^{2}}\right)<1+\varepsilon.$
To sum up,
$\mathfrak{S}_{1}(n)W(z)<318974.199e^{-\gamma}L^{-1}$
then
$\sum_{m\sim U\atop(m,P(z))=1}R(m)<bV^{4}L^{-8},$
where $b=3702014554$.
Naturally, for $n=0$,
$\rho(0)\leq\sum_{m\sim U\atop(m,P(z))=1}R(m)$
and
$\sum_{N/9<\ell\leqslant N}r^{2}(\ell)\leq\rho(0),$
therefore
$\sum_{N/9<\ell\leqslant N}r^{2}(\ell)<bV^{4}L^{-8}.$
We have proved Lemma 5.
4. Proof of the Theorem
In this part, we will prove Theorem.
By the prime number theorem, for
$U=\left(\frac{N}{32(1+\delta)}\right)^{1/4},\quad V=U^{5/6},$
we have
(4.1) $\begin{split}\sum_{N/9<\ell\leqslant N}r(\ell)&\geqslant\sum_{p_{1}\sim U}1\sum_{p_{2}\sim U}1\sum_{p_{3}\sim V}1\sum_{p_{4}\sim V}1\geqslant(1-\varepsilon)\frac{U^{2}V^{2}}{\log^{2}U\log^{2}V}\\ &\geqslant\left(\frac{96}{5}\right)^{2}(1-\varepsilon)U^{2}V^{2}L^{-4}.\end{split}$
then by Cauchy’s inequality and Lemma 5, we have
(4.2) $\begin{split}\left\{\sum_{N/9<\ell\leqslant N}r(\ell)\right\}^{2}&\leqslant\left\{\sum_{\begin{subarray}{c}N/9<\ell\leqslant N\\ r(\ell)>0\end{subarray}}1\right\}\left\{\sum_{N/9<\ell\leqslant N}r^{2}(\ell)\right\}\\ &\leqslant bV^{4}L^{-8}\left\{\sum_{N/9<\ell\leqslant N}1\right\}.\end{split}$
From (4.1), (4.2), we have
$\sum_{N/9<\ell\leqslant N\atop r(\ell)>0}1>\frac{(1-\varepsilon)^{2}}{b}\biggl(\frac{96}{5}\biggr)^{4}U^{4}>\frac{1}{27241.64}N.$
The proof of our Theorem is now complete.
$\square$